Reports 1-1 of 1 Clear search Modify search
AOS (Baffles & Beam dumps)
tomotada.akutsu - 12:15 Friday 03 June 2022 (21009) Print this report
Invac high-power beam dump setting: Day 3

Yano, N. Sato, Akutsu; to follow 20963.


  • Checked the connection of the thermometers on the high power beam dumps to the cryocon; Satou-san might be report later.
  • Installed the Perlier unit onto the Cu flange (Figs. 1 and 2), and connected it with the driver and the power supply, and roughly comfirmed it work. They are all turned off now for safety.


  • Not yet fully completed; we'd like to installed an additional thermometer on the top surface of the optical bench in the IFI chamber, and the design will be done in the next week.
  • We need to discuss within KAGRA how to manage this system. It does not need to keep being turned on unless the invac high-power beam dump would be illuminated by the reflection beam from PRM, especially in high power. It equips a didicated thermometer, and a PID controller, but the discussion point here would be what should be the objective value of the controlled temperature; and maybe also some paramters for the PID.
  • The Peltier unit has a mechanical heat sink and a fan; the fan might be a noise source when it would run, but reminding that this beam dump will be illuminated only PRM would be (intentionally) mis-aligned, meaning the interfeometer would not be in operation with the highest sensitivity, so we may not need to care much such things.
Images attached to this report
Comments to this report:
naohisa.satou - 13:26 Friday 03 June 2022 (21010) Print this report

About wiring for HPBD temperature sensors.

[ Log on 2022.6.1 ]

* Two conversion cables were made and installed to connect the temperature sensor to CRYOCON. These cables will be connected to the signal input terminal of CRYOCON.
* Cable outside the chamber, electrically disconnected from pin 5 of the Dsub cable and the cable shield.

[ Log on 2022.6.3 ]
* We found that there was no need to change the wiring in the chamber, so no changes were made to the wiring in the chamber.
* We connected the cable from the feedthrough to CRYOCON and verified that the temperature is displayed on CRYOCON.
* The correspondence between the display channel of CRYOCON and the installation location of the thermometer is as follows.

[Ch.A] Near the SiC mirror
[Ch.B] Midway point of the heat exhaust structure
[Ch.C] On the optical surface plate (currently not installed)

 Wiring details will be uploaded in JGW-D2214102.

Images attached to this comment
Non-image files attached to this comment
Search Help